# A Survey of Zero-Knowledge Proof Based Verifiable Machine Learning

### v1

@@ Abstract
As machine learning technologies advance rapidly across various domains, concerns over data privacy and model security have grown significantly. These challenges are particularly pronounced when models are trained and deployed on cloud platforms or third-party servers due to the computational resource limitations of users’ end devices. In response, zero-knowledge proof (ZKP) technology has emerged as a promising solution, enabling effective validation of model performance and authenticity in both training and inference processes without disclosing sensitive data. Thus, ZKP ensures the verifiability and security of machine learning models, making it a valuable tool for privacy-preserving AI. Although some research has explored the verifiable machine learning solutions that exploit ZKP, a comprehensive survey and summary of these efforts remain absent. This survey paper aims to bridge this gap by reviewing and analyzing all the existing Zero-Knowledge Machine Learning (ZKML) research from June 2017 to December 2024. We begin by introducing the concept of ZKML and outlining its ZKP algorithmic setups under three key categories: verifiable training, verifiable inference, and verifiable testing. Next, we provide a comprehensive categorization of existing ZKML research within these categories and analyze the works in detail. Furthermore, we explore the implementation challenges faced in this field and discuss the improvement works to address these obstacles. Additionally, we highlight several commercial applications of ZKML technology. Finally, we propose promising directions for future advancements in this domain.
Index Terms:
Zero-knowledge Proof, Machine Learning, Verifiability, Model Security, Data Privacy.
@@ I Introduction
The rapid advancement of artificial intelligence (AI) technologies, epitomized by machine learning (ML), has brought significant transformations to various aspects of human life. Recently, the emergence of generative AI models based on large models has introduced new opportunities in fields such as design and art, software development, publishing, and even finance. However, as model capabilities increase, the demand for computational power in machine learning grows exponentially, necessitating larger datasets and more extensive computational clusters for parallel training. AI business products like ChatGPT [1] and Midjourney [2] present impressive performance, relying not only on sophisticated machine learning models and efficient algorithms but also on substantial investments in data and computational resources for model training.
Due to limitations in computational and data resources, ordinary users such as individuals, small and medium-sized institutions often cannot train ML models locally or even perform inference using trained models. To address this, ML service providers (for example, large companies like Google, Amazon, Alibaba, etc.) offer rental services for ML models, computational resources, and storage space, enabling customers to easily execute ML tasks and integrate them into their applications. For example, ML clients, i.e., individuals, small and medium-sized institutions, utilize Machine Learning as a Service (MLaaS) [3] engines to outsource complex ML models (such as deep neural networks ) training and inference tasks, and performing inferences with trained ML models; the ML service provider receive data from ML clients to execute the ML tasks.
However, the above paradigm raises significant concerns regarding trust and data privacy between ML clients and ML service providers. On one hand, clients are particularly vulnerable due to the sensitive nature of their data, as evidenced by frequent and severe data breaches. Notable examples include the 2022 data breach involving the digital booking and scheduling platform FlexBooker, which resulted in the compromise of personal information for approximately 3.2 million users. Similarly, the theft of a database hosted on Alibaba Cloud exposed the personal information of 1 billion Chinese citizens, along with billions of police records, highlighting the alarming scale and impact of such incidents. Additionally, other issues include ML service providers using underperforming models for inferences while fabricating seemingly flawless results to their clients. As a consequence, clients are often unwilling to provide data containing private information to ML service providers. On the other hand, ML service providers are concerned about their machine learning models being stolen or maliciously compromised. For instance, attackers disguise themselves as clients to implant “backdoors” into the ML service providers’ models, causing it to perform well on normal samples but make specific erroneous inferences on inputs with particular backdoor triggers [4]. Another scenario is that the ML service providers are using a model trained on the wrong data, or using a poorly performing model to make predictions and faking a seemingly perfect result to deceive the client. In summary, due to the trust and data privacy concerns between clients and machine learning service providers, there is a need for privacy-preserving solutions that allow the encrypted data of clients to be used in machine learning tasks while preventing attacks and fraudulent activities on the models of ML service providers, serving trust assurances [5].
Recently, it has been demonstrated that the technique of zero-knowledge proofs (ZKP) can effectively address the aforementioned issues. Consequently, ZKP has been proposed as a solution to implement a verifiable machine learning framework, known as zero-knowledge machine learning (ZKML). ZKP is a type of cryptographic technique that allows one party to prove the truth of a statement to another party without revealing any information beyond the statement itself [6]. They have the potential to effectively solve the privacy protection and tissues of data and models in MLaaS, as well as the trustworthiness of model computation results. Within the framework of ZKML, the correctness of data and the correctness of model parameters or execution results can be considered as a ”statement” that the ZKP is used to prove. For example, in a ZKP-based training task of verifiable machine learning, the client needs to prove to the ML service provider that the training data provided by the client is indeed correct; conversely, the ML service provider needs to prove to the client that the model parameters (or inference results) are indeed obtained by training (or performing inference) on the dataset received from the client. ZKP allows the two parties to trust the statements without learning the details of task executions.
Fig. 1 illustrates one of ZKML setups for executing inference tasks in ZKML. In this ZKML setup, the machine learning service provider acts as the prover in the ZKP algorithm, and the client acts as the verifier. The function $F(\ast)$ represents the relationship the ZKP algorithm wants to prove. The input data for the ZKP algorithm is divided into public input (known to both the prover and verifier) and the private witness (known only to the prover). To execute this ZKML setup, a client first sends their data $x$ to the ML service provider as the data input to the ML model; and then the ML service provider supplies the model parameters $w$ together with the data $x$ to completes the machine learning inference task by computing the model inference result $r=W(x,w)$, where $W(\ast)$ denotes the model inference computation; after that, the machine learning service provider runs the proving function of ZKP to generate a zero-knowledge proof $\pi$ for that the relationship $F(x,r,w)=W(x,w)-r=0$ is hold, where $(x,r)$ are set as the public input of the ZKP algorithm and $w$ is set to as the private witness of the ZKP algorithm; finally, the client runs the verifying function of ZKP to verify the proof $\pi$ and accepts the inference result $r$ if the proof $\pi$ is verified successfully. Through this ZKML setup, the client can trust that the inference result $r=W(x,w)$ provided by the ML service provider is indeed obtained by executing the machine learning model on their provided data $x$, even though they do not learn the specific model parameters $w$ and the exact inference computation process of $W(\ast)$. In this way, a form of “trustworthy” machine learning is achieved via the verifiability of ZKP.
Figure 1: An illustrating setup of ZKML Framework.
Thanks to the growing interest in privacy-preserving and trust ML technologies and the advancements in zero-knowledge proof techniques, there are now numerous works on ZKML, and new proposals continue to emerge. ZKML is a novel verifiable ML framework at the intersection of machine learning and cryptography, characterized by a diverse and complex research landscape. Therefore, systematically reviewing its development trajectory and summarizing current progress is crucial for advancing future research. Currently, there are limited survey studies on ZKML research works and no comprehensive coverage about the latest research up to now, lacking depth understanding about the relationship between research works. The existing survey works about ZKML are [7, 8, 9], and we summarized them as follows. In [7], Modulus Lab conducted inductive and comparative studies on the verifiability of machine learning inference task based on ZKP, confirming the feasibility of ZKP in verifiable machine learning. Using multilayer perceptrons (MLP) as the benchmark machine learning model, they compared the performances of six different ZKP systems, including Groth16, Gemini, Winterfell, Halo2, Plonky2, and zkCNN [10], in terms of proof time and memory consumption during machine learning inferences. Additionally, through comprehensive experimental validation, this work explained the impact of different ZKP systems on the efficiency of machine learning verifiable inference works. However, their analysis of how ZKML works enhance efficiency is relatively limited (all the works mentioned in [7] are also included for investigations in this survey). Sathe et al. [8] introduced and analyzed ZKML works up to 2023, providing detailed descriptions of zkCNN [10], ezDPS [11], Xing [12], and Mystique [13]. However, this survey covers fewer works and does not provide a comparative investigation of the existing works; the authors of [8], individually described the specific contents of the four works mentioned, failing to comprehensively present the latest developments and directions in the field. Recently, Xing et al. [9] provided a comprehensive survey of ZKML, including definitions, properties, and challenges. This survey covered all relevant ZKML works up to June 2023. The existing works discussed in this survey were divided into two application categories and further classified based on technical characteristics, providing researchers with a more comprehensive reference.
To bridge the gap and enhance existing reviews and surveys on ZKML, we conducted a comprehensive investigation summarizing ZKML works from June 2017 to December 2024. The main contributions of this survey paper are as follows:
  • Systematic Investigation of ZKML Research: This paper offers a structured examination of ZKML works, organizing, classifying, and summarizing nearly all significant contributions from June 2017 to December 2024, encompassing a total of 27 notable studies.
  • Categorization and Discussion of Technical Improvements: We categorize the technical improvement works that aim to address the implementation challenges of ZKML research into two primary dimensions (i.e., improving the generality and efficiency of ZKML) and outline the evolutionary process of these advancements.
  • Introduction of Commercial Applications: Moving beyond academic contexts, this paper explores commercial applications of ZKML, demonstrating its relevance and potential impact in industry.
  • Future Directions and Technical Challenges: By analyzing the current state of ZKML research and identifying key technical challenges, this paper proposes potential avenues for future development, offering guidance and inspiration for subsequent researchers.
@@
The remainder of the paper is organized as follows: Section II introduces the background of machine learning and zero-knowledge proofs, along with three categories of verifiable machine learning. Section III presents existing research works for verifiable machine learning based on zero-knowledge proofs and their development process. Section IV discusses commercial applications of ZKML. Section V not only summarizes the paper, but also proposes future directions for ZKML development. Table I lists the main abbreviations used in the paper and their definitions.
TABLE I: Abbreviations Table
Abbreviations Full Name
@@
zk-SNARK Zero-Knowledge Succinct
Non-interactive Argument of Knowledge
@@
PCA principal components analysis
@@ II Background
This section starts by providing the technical background of machine learning and zero-knowledge proofs. It then delves into the details of verifiable machine learning and concludes with a comparison of the strengths and limitations of zero-knowledge proofs relative to other security techniques in the context of verifiable machine learning.
@@ II-A Machine Learning
Machine Learning (ML) [14], a subfield of AI, enables the development of models from diverse types of data, such as numerical values, textual content, images, and user interactions. These ML models are applied to tasks like pattern recognition, problem-solving, and making predictions [15]. ML encompasses various approaches, including supervised learning, unsupervised learning, reinforcement learning, and more. In this survey, we primarily focus on supervised learning, as it is the most common form of ML and serves as the foundation for the most of existing zero-knowledge-based verifiable machine learning research.
In supervised learning, the goal is to train a model that can capture the mapping $g(\cdot)$ from inputs $X$ (also referred to as data features) to outputs $Y$ (also known as labels), expressed as $Y=g(X)$. We denote a ML model by $f_{\theta}(\cdot)$, where $\theta$ represents the parameters of the model. The model is trained using a training algorithm and a dataset $D_{train}=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})\}$, where each pair $(X_{i},Y_{i})$ represents an input and its corresponding output label. In the training process, the parameters of the model are optimized by minimizing the total loss over the training dataset, formalized as: $\theta=\arg\mathop{\min}\limits_{\theta^{\prime}}\sum\nolimits_{\left({{Y_{i}},{X_{i}}}\right)\in{D_{train}}}{l\left({{Y_{i}},{f_{\theta^{\prime}}}\left({{X_{i}}}\right)}\right)}$, where ${l\left({\cdot,\cdot}\right)}$ denotes the chosen loss function. Once the model $f_{\theta}(\cdot)$ is trained, it can be used to infer the output label $Y^{\prime}$ for new input data $X$ that were not part of the training dataset, expressed as $Y^{\prime}=f_{\theta}(X)$. The goal is for the learned model $f_{\theta}(\cdot)$ to approximate the true mapping $g(\cdot)$ as closely as possible, ensuring that the predicted output label $Y^{\prime}$ closely matches the true label $Y$.
Various models are used to represent the mapping between input data features and output labels, including linear regression, decision trees, support vector machines, and deep neural networks. Each model has distinct characteristics and is suited to specific scenarios. Given the popularity and effectiveness of deep neural networks in modern applications, as well as the fact that nearly all existing work on ZKML focuses on deep neural networks, we adopt them as the default ML model in this survey unless otherwise specified. Consequently, the model parameters refer to the weights of the connections between the neuron units across the layers of the deep neural network.
@@ II-B Zero-Knowledge Proof
Zero-knowledge proof (ZKP) is a cryptographic technique that proves a statement’s validity without revealing any private information about the statement. There are several types of zero-knowledge proof systems [16], for example, zero-knowledge succinct non-interactive argument of knowledge (zk-SNARK) [17], zero-knowledge scalable transparent argument of knowledge (zk-STARK) [18], and others.
@@
The computation of a ZKP system is usually represented by an arithmetic circuit that consists of the basic arithmetic operations of addition, subtraction, multiplication, and division. [^1] Actually, the ZKP systems of zk-SNARK are always associated with arithmetic circuit representations; zk-STARK can theoretically support arithmetic circuit representations, but it typically uses the more efficient arithmetic intermediate representation (AIR) [19] to construct ZKP systems. An $\mathbb{F}$-arithmetic circuit is a circuit in which all inputs and all outputs are elements in a field $\mathbb{F}$. Consider an $\mathbb{F}$-arithmetic circuit $C$ that has an input $x\in\mathbb{F}^{n}$, an auxiliary input $W\in\mathbb{F}^{h}$ called a witness, and an output $C(x,W)\in\mathbb{F}^{l}$, where $n,h,l$ are the dimensions of the input, auxiliary input, and output, respectively. The arithmetic circuit satisfiability problem of the $\mathbb{F}$-arithmetic circuit $C$ is captured by the relation: $\mathcal{R}_{C}=\{(x,W)\in\mathbb{F}^{n}\times\mathbb{F}^{h}:C(x,W)=0^{l}\}$ and its expression is $\mathcal{L}_{C}=\{x\in\mathbb{F}^{n}:\exists W\in\mathbb{F}^{h}\text{ s.t. }C(x,W)=0^{l}\}$. A ZKP system consists of three algorithmic components [17]:
  • $\blacksquare$
@@
    $(PK,VK)\leftarrow KEYGEN(1^{\lambda},C)$ is the key generation algorithm that generates the proving key $PK$ and the verification key $VK$ by using a predefined security parameter $\lambda$ and an arithmetic circuit representation $C$.
  • $\blacksquare$
@@
    $\pi\leftarrow PROVE(PK,x,W)$ is the proof generation algorithm that generates a proof $\pi$ based on the proving key $PK$, the input $x$ and the witness $W$.
  • $\blacksquare$
@@
    $1/0\leftarrow VERIFY(VK,x,\pi)$ is the proof verification algorithm that outputs a decision to accept or reject $\pi$ using $VK$, $x$ and $\pi$ as the input.
The proving key PK and the verification key VK generated by the KEYGEN algorithm are treated as the public parameters. The PROVE algorithm is executed by the prover, and the VERIFY algorithm is executed by the verifier. Witness W is the secret owned by the prover that they do not want to reveal to others and yet wants to prove that they know the secret.
Different ZKP systems are computationally suited to different types of arithmetic circuits, each with distinct characteristics. As a result, various ZKP systems excel in different application domains. For instance, zk-SNARKs are particularly effective at handling arithmetic circuits represented in the Rank-1 Constraint System (R1CS) form; whereas ZKP systems based on sum-check [20] and GKR [21] protocols are more suitable for dealing with arithmetic circuits that exhibit a layered structure. This layered structure of arithmetic circuits is particularly relevant in the context of deep neural network models in machine learning, which typically possess a layered structure. Consequently, many ZKML works targeting deep neural networks leverage sum-check and GKR protocols for enhanced efficiency and performance.
In general, the following two technical advantages of ZKP systems are useful for verifiable machine learning and thus are exploited by ZKML. The VERIFY algorithm can be executed significantly faster than the PROVE algorithm. The VERIFY algorithm does not require access to the private witness. The first advantage enables machine learning clients to save computational resources compared to performing all computations themselves. The second advantage ensures that sensitive information remains protected throughout the entire machine learning process, preventing any disclosure.

### v2

@@ Abstract
Machine learning is increasingly deployed through outsourced and cloud-based pipelines, which improve accessibility but also raise concerns about computational integrity, data privacy, and model confidentiality. Zero-knowledge proofs (ZKPs) provide a compelling foundation for verifiable machine learning because they allow one party to certify that a training, testing, or inference result was produced by the claimed computation without revealing sensitive data or proprietary model parameters. Despite rapid progress in zero-knowledge machine learning (ZKML), the literature remains fragmented across different cryptographic settings, ML tasks, and system objectives. This survey presents a comprehensive review of ZKML research published from June 2017 to August 2025. We first introduce the basic ZKP formulations underlying ZKML and organize existing studies into three core tasks: verifiable training, verifiable testing, and verifiable inference. We then synthesize representative systems, compare their design choices, and analyze the main implementation bottlenecks, including limited circuit expressiveness, high proving cost, and deployment complexity. In addition, we summarize major techniques for improving generality and efficiency, review emerging commercial efforts, and discuss promising future directions. By consolidating the design space of ZKML, this survey aims to provide a structured reference for researchers and practitioners working on trustworthy and privacy-preserving machine learning.
@@ I Introduction
Artificial intelligence (AI), with machine learning (ML) as its core engine, is increasingly embedded in modern applications such as content generation, software engineering, finance, healthcare, and intelligent decision support. This progress is driven by rapid advances in model architectures, optimization methods, and training infrastructure. At the same time, the computational cost of state-of-the-art ML continues to rise sharply: larger models require larger datasets, more memory, and increasingly specialized hardware. Commercial systems such as ChatGPT [1] and Midjourney [2] exemplify this trend, as their performance depends not only on algorithmic innovations but also on large-scale investments in data and compute resources.
This resource asymmetry has accelerated the adoption of Machine Learning as a Service (MLaaS) [3]. In the MLaaS paradigm, clients outsource training, testing, or inference to external service providers that host the required models, hardware, or data-processing pipelines. While this model improves accessibility, it also creates two intertwined concerns. First, clients may be unwilling to expose sensitive data to external providers because of privacy risks and frequent real-world data breaches. Second, service providers often regard trained models as valuable intellectual property and are reluctant to reveal model parameters or internal execution details. At the same time, clients still need assurance that the returned model, metric, or prediction was produced by the declared computation rather than fabricated, shortcut, or generated using an unintended model or dataset. These tensions make trust and privacy central challenges in outsourced ML [4, 5].
These requirements motivate verifiable machine learning, whose goal is to establish both computational integrity and privacy. Zero-knowledge proof (ZKP) is particularly well suited to this objective. A ZKP allows one party to convince another party that a statement is true without revealing any information beyond the validity of that statement [6]. In the ZKML setting, the statement can encode claims such as “this prediction equals the output of the committed model on the declared input” or “these updated parameters were produced by the declared training procedure on the committed dataset.” By coupling ML computation with cryptographic proof generation, zero-knowledge machine learning (ZKML) enables mutually distrustful parties to establish trust in an outsourced ML task without disclosing sensitive training data, inference inputs, or proprietary model parameters.
[…]
I-A Illustrative ZKML Inference Workflow
[…]
I-B Positioning Against Existing Surveys
Interest in ZKML has grown rapidly in both academia and industry, and the literature now spans circuit design, proof-system engineering, framework construction, and application-specific optimization. However, survey coverage of the area is still limited and fragmented. Existing surveys usually focus on only one slice of the field. Modulus Labs [7] benchmarked practical ZKP systems for ML inference and highlighted proof-time and memory trade-offs. Sathe et al. [8] summarized several representative early ZKML systems but covered only a relatively small set of works. Xing et al. [9] focused on decentralized ML in communication networks and covered the literature up to June 2023. The open-source review in [10] examined practical ZKP frameworks and development environments, but its main emphasis is framework usability rather than the broader ZKML research landscape. Therefore, the field still lacks an up-to-date survey that simultaneously covers algorithmic categories, technical evolution, implementation bottlenecks, commercial deployment, and future hybrid cryptographic directions. Table I compares these existing efforts with our survey.
TABLE I: Comparison of existing ZKML surveys and this survey.
Aspect Ref. Main Contributions
Benchmarking practical ZKP systems for ML inference [7] - Benchmarking six representative ZKP systems for ML inference; - Highlighting proof-time and memory trade-offs across different constructions; - Discussing practical deployment challenges through real-world case studies.
Early survey of representative ZKML systems [8] - Summarizing early representative systems such as zkCNN, ezDPS, Xing, and Mystique; - Discussing CNN-oriented verification optimizations and implementation ideas.
ZKP-based decentralized ML in communication networks [9] - Formalizing a four-algorithm ZKP-VML model; - Proposing distributed verification architectures (1-to-1, 1-to-$m$, and $m$-to-1); - Analyzing attack surfaces and defense mechanisms in networked settings.
Assessment of open-source ZKP frameworks [10] - Comparing open-source ZKP frameworks, including constructions, environments, and benchmarks; - Evaluating scalability, runtime, proof size, and usability; - Providing practical recommendations for framework selection.
Our work - Covering ZKP-based verifiable training, testing, and inference in a unified framework; - Reviewing the literature from 2017 to 2025 and synthesizing the technical evolution of the field; - Examining commercial applications and hybrid cryptographic frameworks for practical deployment; - Identifying open challenges and future directions in efficiency, generality, privacy, and scalability.
I-C Contributions
To address the above gap, we conducted a systematic survey of ZKML studies published from June 2017 to August 2025. We collected academic papers from major scientific databases, including ACM Digital Library, IEEE Xplore, Elsevier, MDPI, and SpringerLink. Since ZKML has also been actively explored in industry, we additionally reviewed white papers and technical blogs. Using keyword combinations related to zero-knowledge proofs and machine learning, we constructed a literature corpus that captures both foundational research and recent implementation-oriented developments. Based on this corpus, the main contributions of this survey are as follows:
  • Comprehensive Research Landscape: We provide a structured examination of ZKML research, organizing and analyzing 27 representative studies from June 2017 to August 2025.
  • Technical Evolution and Implementation Insights: We categorize the major technical improvements in ZKML, especially those targeting efficiency and generality, and explain the design trade-offs behind these advances.
  • Bridge from Research to Practice: We review commercial applications and practical frameworks for ZKML, highlighting how the field is evolving from conceptual prototypes toward deployable systems.
  • Open Problems and Future Directions: We identify key technical bottlenecks and promising research directions, including computational optimization, model support, privacy enhancement, and scenario expansion.
@@ I-C Contributions
The remainder of this paper is organized as follows. Section II reviews the background of machine learning, zero-knowledge proofs, verifiable machine learning, and related security techniques. Section III surveys representative ZKML studies and their technical evolution. Section IV discusses commercial applications of ZKML. Section V examines hybrid cryptographic frameworks for verifiable ML. Section VI outlines future directions for ZKML, and Section VII concludes the paper. Table II lists the main abbreviations used throughout this paper.
TABLE II: Main abbreviations used throughout this paper.
Abbreviation Full Name
@@
zk-SNARK Zero-Knowledge Succinct Non-interactive Argument of Knowledge
@@
PCA Principal Component Analysis
@@ II Background
This section reviews the technical foundations needed for the remainder of the survey. We first introduce the basic notation of machine learning and zero-knowledge proofs, then define the three main categories of verifiable machine learning, and finally compare ZKP with other representative security techniques used in trustworthy ML systems.
@@ II-A Machine Learning
Machine learning (ML) [11] is a subfield of AI that learns predictive or decision-making functions from data and applies them to tasks such as classification, regression, pattern recognition, and control [12]. Major paradigms include supervised learning, unsupervised learning, reinforcement learning, and self-supervised learning. Since most existing ZKML systems focus on proving the correctness of supervised training, testing, or inference, we adopt supervised learning as the default setting in this survey.
In supervised learning, let $\mathcal{X}$ and $\mathcal{Y}$ denote the input space and label space, respectively, and let $g:\mathcal{X}\rightarrow\mathcal{Y}$ denote the target mapping. We represent an ML model by $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\theta$ denotes its trainable parameters. The model is trained on a dataset $\mathcal{D}_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where $x_{i}\in\mathcal{X}$ is an input sample and $y_{i}\in\mathcal{Y}$ is its corresponding label. The training procedure optimizes the model parameters by minimizing the empirical loss over $\mathcal{D}_{\mathrm{train}}$, namely, $\theta=\arg\min_{\theta^{\prime}}\sum_{i=1}^{n}\ell\bigl(f_{\theta^{\prime}}(x_{i}),y_{i}\bigr)$, where $\ell(\cdot,\cdot)$ denotes the loss function. After training, for a new input $x\in\mathcal{X}$, the model predicts the output $\hat{y}=f_{\theta}(x)$. The objective is for the learned model $f_{\theta}$ to approximate the target mapping $g$, so that the prediction $\hat{y}$ is as close as possible to the true label $y$.
Various models can be used to represent the mapping from input features to output labels, including linear regression, decision trees, support vector machines, and deep neural networks. Each model has distinct characteristics and is suited to specific scenarios. Given the popularity and effectiveness of deep neural networks in modern applications, as well as the fact that nearly all existing work on ZKML focuses on deep neural networks, we adopt deep neural networks as the default ML model in this survey unless otherwise specified. Accordingly, $\theta$ denotes the trainable parameters of the network, including weights and biases.
@@ II-B Zero-Knowledge Proof
Zero-knowledge proof (ZKP) is a cryptographic technique that enables a prover $\mathcal{P}$ to convince a verifier $\mathcal{V}$ that a statement is true without revealing any additional information beyond the validity of that statement. Representative ZKP families include zero-knowledge succinct non-interactive arguments of knowledge (zk-SNARKs) [13], zero-knowledge scalable transparent arguments of knowledge (zk-STARKs) [14], and related constructions [15].
@@
For many practical ZKP systems, the statement to be proven is encoded as an arithmetic circuit composed of addition, subtraction, multiplication, and division operations.111zk-SNARK systems are typically associated with arithmetic-circuit representations; zk-STARK can also support arithmetic-circuit representations, but in practice it often uses the more efficient arithmetic intermediate representation (AIR) [16]. An $\mathbb{F}$-arithmetic circuit is a circuit in which all inputs and outputs are elements of a field $\mathbb{F}$. Consider an $\mathbb{F}$-arithmetic circuit $C$ with public input $x\in\mathbb{F}^{n}$, witness $w\in\mathbb{F}^{h}$, and output $C(x,w)\in\mathbb{F}^{l}$, where $n$, $h$, and $l$ denote the dimensions of the public input, witness, and output, respectively. The corresponding arithmetic-circuit satisfiability relation is $\mathcal{R}_{C}=\{(x,w)\in\mathbb{F}^{n}\times\mathbb{F}^{h}:C(x,w)=0^{l}\}$, and its induced language is $\mathcal{L}_{C}=\{x\in\mathbb{F}^{n}:\exists w\in\mathbb{F}^{h}\text{ s.t. }C(x,w)=0^{l}\}$. A ZKP system then consists of the following three algorithmic components [13]:
@@
  • $(\mathsf{PK},\mathsf{VK})\leftarrow\mathsf{KEYGEN}(1^{\lambda},C)$ is the key-generation algorithm that outputs the proving key $\mathsf{PK}$ and the verification key $\mathsf{VK}$ using the security parameter $\lambda$ and the arithmetic circuit $C$.
@@
  • $\pi\leftarrow\mathsf{PROVE}(\mathsf{PK},x,w)$ is the proof-generation algorithm that outputs a proof $\pi$ given the proving key $\mathsf{PK}$, the public input $x$, and the witness $w$.
@@
  • $1/0\leftarrow\mathsf{VERIFY}(\mathsf{VK},x,\pi)$ is the proof-verification algorithm that outputs an accept/reject decision using $\mathsf{VK}$, $x$, and $\pi$ as input.
The proving key $\mathsf{PK}$ and the verification key $\mathsf{VK}$ generated by $\mathsf{KEYGEN}$ are public parameters. The prover $\mathcal{P}$ executes $\mathsf{PROVE}$, and the verifier $\mathcal{V}$ executes $\mathsf{VERIFY}$. The witness $w$ is the secret held by $\mathcal{P}$ that should remain hidden while still being proven known.
Different ZKP systems are suited to different circuit representations and computational structures. For example, zk-SNARKs are particularly effective for arithmetic circuits expressed in Rank-1 Constraint System (R1CS) form, whereas systems based on the sum-check [17] and GKR [18] protocols are often more suitable for layered computations. This distinction is highly relevant to ML, because deep neural networks naturally exhibit layered structures. Consequently, many ZKML systems for neural-network verification build on sum-check- or GKR-style proving frameworks to improve efficiency.
For ZKML, two properties of ZKP are particularly important. First, $\mathsf{VERIFY}$ is usually much cheaper to execute than $\mathsf{PROVE}$, which allows a resource-constrained client to validate an expensive outsourced ML computation efficiently. Second, $\mathsf{VERIFY}$ does not require access to the private witness, which means that sensitive information such as model parameters, training data, or internal intermediate states can remain hidden throughout the verification process.